VEGAS Pulsar Project Meeting: 2015 May 4th 2:00 - 3:00pm ET

Room / Connection Details

  • GB-137 / Soc 280
  • 192.33.117.12##7144
  • 434-817-6443

Housekeeping Details

  • Main wiki page
  • gbsapp Mailing List
  • Location and regular date and time of meeting - agreed weekly on Mondays; GB-137 / Soc-280; 2-3pm ET
  • ETK codes:
    • Electronics: 432510.GB3015
    • Software: 893252.GB3015
    • Science: 432540.GB3015

Resource Allocations

GUPITA, ARINDAM 1.00
FORD, JOHN 0.10
MCCULLOUGH, RANDY 0.40
RAY, JASON 0.55
BLOSS, MARTY 0.05
CREAGER, RAY 0.8
RICHMOND DECKER, JUSTIN 0.0
LYNCH, RYAN 0.25
DEMOREST, PAUL 0.25
PRESTAGE, RICHARD 0.1

Present

  • Marty, Paul D., Jason, Randy, Arindam, Richard, Justin, Scott, Ryan

Agenda

General Updates

  • New BOF files
    • Status and summary page
    • Device name rationalization update
      • Arindam put together a detailed list
      • The initial plan has been updated after feedback from Joe and others and is now available here [Updated 05-11-15 ASG]
    • Currently testing Xilinx v14.7
      • Testing ADC blocks by checking for recovering of injected tone * 1500 MHz modes look OK8 * 800 MHz modes show elevated noise floor and broader spectral peak, possible due to clock instability? * Will proceed to build all bofs under v14.7 once ADC tests are completed
        • spectrum 147 800-01.png
      • ADC test results [05-11-15 ASG]:
        • Issue with the 800 MHz ADC resolved. Both 1500 MHz and 800 MHz built on Xilinx 14.7 now seem to work fine.
        • spectrum 147 800 updated.png
        • spectrum 147 1500-01.png
  • Manager
    • guppi_daq_server is still timing out when switching between certain modes
      • Seems to occur when changing number of channels, but not when switching between incoherent and coherent modes with same nchan
      • UPDATE: Justin made some changes to the manager and it now seems to work when switching between different incoherent modes, but coherent modes changes still timing out
      • Justin will compile a list of mode changes that trigger this behavior

Discussion

  • Periodicities in TOAs tracked down to incorrect/unexpected values in valon synthesizer
    • Reset seems to have fixed this issue
    • Need to make sure that initial configuration is being done correctly
  • Paul believes a more precise test set up for the valons can be achieved by injecting a pure tone and using a coherent mode to recover it
    • Jason will set this up and Ryan will perform tests

Date / Time of Next Meeting

  • May 11

-- RyanLynch - 2015-05-04
Topic attachments
I Attachment Action Size Date Who Comment
spectrum_147_1500-01.pngpng spectrum_147_1500-01.png manage 152 K 2015-05-11 - 13:15 ArindamSengupta 1500 MHz ADC built on Xilinx 14.7
spectrum_147_800-01.pngpng spectrum_147_800-01.png manage 142 K 2015-05-11 - 13:28 ArindamSengupta 800 MHz ADC (bad spectrum)
spectrum_147_800_updated.pngpng spectrum_147_800_updated.png manage 131 K 2015-05-11 - 13:14 ArindamSengupta 800 MHz ADC built on Xilinx 14.7
Topic revision: r5 - 2015-05-11, ArindamSengupta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback